# DBConformer: Dual-Branch Convolutional Transformer for EEG Decoding

### v1

@@ Abstract
Electroencephalography (EEG)-based brain-computer interfaces (BCIs) transform spontaneous/evoked neural activity into control commands for external communication. While convolutional neural networks (CNNs) remain the mainstream backbone for EEG decoding, their inherently short receptive field makes it difficult to capture long-range temporal dependencies and global inter-channel relationships. Recent CNN-Transformer (Conformers) hybrids partially address this issue, but most adopt a serial design, resulting in suboptimal integration of local and global features, and often overlook explicit channel-wise modeling. To address these limitations, we propose DBConformer, a dual-branch convolutional Transformer network tailored for EEG decoding. It integrates a temporal Conformer to model long-range temporal dependencies and a spatial Conformer to extract inter-channel interactions, capturing both temporal dynamics and spatial patterns in EEG signals. A lightweight channel attention module further refines spatial representations by assigning data-driven importance to EEG channels. Extensive experiments on five motor imagery (MI) datasets and two seizure detection datasets under three evaluation settings demonstrate that DBConformer consistently outperforms 10 competitive baseline models, with over 8$\times$ fewer parameters than the high-capacity EEG Conformer baseline. Further, the visualization results confirm that the features extracted by DBConformer are physiologically interpretable and aligned with sensorimotor priors in MI. The superior performance and interpretability of DBConformer make it reliable for robust and explainable EEG decoding. Code is publicized at https://github.com/wzwvv/DBConformer.
Index Terms:
Brain-computer interface, electroencephalography, motor imagery, seizure detection, convolutional neural networks, Transformer
@@ I Introduction
A brain-computer interface (BCI) provides a direct communication pathway between cortical activity and external devices [1]. Non-invasive BCIs based on electroencephalography (EEG) have gained prominence because EEG sensors are inexpensive, portable, and safe for extended use [2]. Nevertheless, the low signal-to-noise ratio, non-stationarity, and significant inter-subject variability of scalp EEG impose stringent requirements on the decoding model that translates raw signals into task-relevant information. This study focuses on two significant EEG decoding tasks: motor imagery (MI) classification [3], which enables hands-free control of prostheses and wheelchairs, and epileptic seizure detection [4], a task of automated neurological monitoring.
During MI, users imagine the movement of specific body parts (e.g., left hand, right hand, both feet, or tongue), producing event-related desynchronization/synchronization patterns over the brain’s motor cortex [5]. Despite substantial progress, decoding MI from EEG signals remains challenging.
Epilepsy, affecting over 70 million people worldwide [6], is characterized by recurrent seizures. Continuous EEG monitoring remains the gold standard for seizure detection, but manual annotation by doctors is labor-intensive and costly. Automatic seizure detection must cope with heterogeneous seizure morphologies, varying recordings, and scarce labeled data [7].
Decoding models are essential for effective EEG decoding and has been a research hotspot of BCIs. Three trends have emerged in the design of EEG decoding models:
@@
    Model Type: Classic models based on convolutional neural networks (CNNs), such as EEGNet [8], SCNN, and DCNN [9], exploit local spatio-temporal patterns but struggle with global temporal dependencies. CNN-Transformer hybrids (Conformers) extend CNN locality with self-attention based global context. Yet, prior work [10, 11, 12, 13] largely stacks the two modules serially, restricting interaction between local and global representations and neglecting explicit spatial modeling.
@@
Overall, the latest serial Conformer hybrids first compress EEG trials into patch embeddings by CNN blocks and then utilize self-attention to capture global long-range dependencies. While effective for certain tasks, this cascade forces all spatial information to pass through a temporal bottleneck, preventing interaction between local and global characteristics. Notably, some Conformer hybrids have large model sizes, e.g., EEG Conformer contains around 0.8M parameters, which increases computational cost and may overfit on limited EEG data. Besides, spatial inter-channel relationships, often captured in early CNN layers, may be overwhelmed in later blocks, limiting the model’s capacity to leverage spatial information crucial for MI classification and seizure detection. A balanced, parallel design that treats temporal and spatial patterns symmetrically remains absent.
To address the above limitations, we propose DBConformer, a dual-branch convolutional Transformer architecture that performs parallel spatio-temporal representation learning for EEG decoding. The temporal Conformer (T-Conformer) extracts hierarchical temporal dynamics by coupling depth-wise temporal convolutions with multi-head self-attention, whereas the spatial Conformer (S-Conformer) leverages point-wise convolutions and channel-wise self-attention to model spatial dependencies. A lightweight channel attention module adaptively refines spatial saliency before branch fusion. In each branch, CNN extracts temporal/spatial local representations, and Transformer captures global dependencies in the temporal or channel horizon. The comparison of network architectures among CNNs, traditional serial Conformers, and the proposed DBConformer, illustrated in Fig. 1.
Figure 1: Comparison of network architectures among CNNs (EEGNet[8], SCNN [9], DCNN [9], etc), traditional serial Conformers (EEG Conformer [10], CTNet [11], etc), and the proposed DBConformer. DBConformer has two branches that parallel capture temporal and spatial characteristics.
@@
    Extensive experiments on five MI datasets and two seizure detection datasets validated that DBConformer consistently outperforms 10 competitive baseline models. Furthermore, it achieves this superior performance with over 8$\times$ fewer parameters than the high-capacity EEG Conformer baseline, demonstrating both effectiveness and efficiency.
@@ II Related Works
FBCNet [14] Parallel 1 Spectral MI
@@ II-B Serial CNN-Transformer Hybrids
To incorporate global information, recent models augment CNNs with Transformers, typically in a serial architecture. For example, EEG Conformer [10], CTNet [11], EEG-Deformer [12], and SE-TSS-Transformer [13] compress EEG trials into patch embeddings using convolutional blocks and subsequently apply Transformer layers to capture long-range dependencies. However, this serial design constrains information flow to a single pathway, where local and global features are processed sequentially. As a result, spatial inter-channel relationships, often captured in early CNN layers, may be overwhelmed in later blocks, limiting the model’s capacity to leverage spatial information.
@@ II-C Other Architectures
Some research explores architectures beyond CNNs, including full self-attention mechanisms [33] and graph neural networks [34, 35, 36]. Transformer-based models capture long-range dependencies, and graph neural networks leverage electrode topology to refine spatial characteristics. These models are typically parameter-intensive and rarely integrate local feature extraction with global modeling.
@@ III-B Data Normalization
The raw EEG trials of a batch are $\mathbf{X}\!\in\!\mathbb{R}^{B\times C\times T}$, where $B$ is the batch size, $C$ the number of channels, and $T$ the number of time samples. We apply Euclidean alignment (EA) [37] to align the original trials from each subject in the Euclidean space. Assume a subject has $n$ trials, EA can be formulated as:
@@ III-C Network Architecture
T-Conformer Separable Conv1D $F$ (1,)(1,)( 1 , ) $C\cdot F$ $(B,F,T)$ Spatial filtering
@@
[…]
@@
III-C1 T-Conformer
@@
Inspired by prior CNN backbones in EEG decoding [15], we design the module with a depthwise separable 1D convolution, a 1D temporal convolution, and an average pooling layer. The first separable convolution has $F$ kernels with a stride of (1,)(1,)( 1 , ), followed with batch normalization. The second temporal convolution applies $F$ filters, accompanied by batch normalization, a GELU activation, and a dropout with a probability $p=0.5$. Then, an average pooling with kernel size $(1,W)$ is applied along the time dimension to form non-overlapping temporal patches $\mathbf{Z}_{\text{t}}$. Finally, the patches are reshaped via transposition. The number of filters $F$ is set equal to the Transformer embedding dimension $D$, thereby eliminating the need for a separate linear projection.
@@
$\mathbf{F}_{\text{t}}=\frac{1}{P}\sum_{i=1}^{P}\mathbf{Z}_{t}^{\text{out}}[:,i,:]\in\mathbb{R}^{B\times D},$ (5)
@@ III-C2 S-Conformer
Figure 3: Details of the proposed channel attention module.
@@
Then, each token is transformed into attention-weighted sum across channels:
@@
By repeating across all $B$ samples in a batch, the final batch-level spatial representation $\mathbf{F}_{\text{s}}\in\mathbb{R}^{B\times D}$ can be obtained.
III-C3 Feature Fusion and Classification
@@
which is then passed through a multi-layer perceptron classifier to generate the final predictions $\hat{y}_{i}$. The classifier consists of three fully connected layers with intermediate ELU activations and dropout regularization.
@@ IV Experiments and Results
This section details the datasets, experiments, and analyses. We implemented and fairly evaluated ten state-of-the-art EEG decoding models, including CNN-based, CNN-Transformer hybrid, and CNN-Mamba hybrid. Code for DBConformer, along with all compared baseline models, is available on GitHub [^1] https://github.com/wzwvv/DBConformer, serving as a benchmark codebase for EEG decoding.
@@ IV-A1 MI Datasets
We adopted five publicly available MI datasets from the MOABB benchmark [39] and BCI competitions:
@@
  • Zhou2016 [42]: Comprises EEG trials from 4 subjects performing left/right hand MI, recorded with 14 channels at 250 Hz. The first session was used in the experiments.
  • Blankertz2007 [43]: From BCI Competition IV dataset 1, includes 7 subjects with 59-channel EEG sampled at 250 Hz. Subjects 1 and 5 performed left hand/right foot MI, and the other five subjects performed left/right hand MI.
  • BNCI2014002 [44]: Includes 14 subjects, each performing eight runs right hand/feet MI, using 15 EEG channels recorded at 512 Hz. The first five training runs were used in the experiments.
@@ IV-A2 Seizure Detection Datasets
TABLE III: Summary of the five MI datasets and two seizure datasets.
@@
IV-A3 Data Preprocessing
For BNCI2014001, BNCI2014004, Zhou2016, and BNCI2014002 datasets, the standard preprocessing steps in MOABB, including notch filtering, band-pass filtering, etc., were performed to ensure the reproducibility. For the Blankertz2007 dataset, EEG trials were first band-pass filtered between 8 and 30 Hz. Then, trials between [0.5, 3.5] seconds after the cue onset were segmented and downsampled to 250 Hz.
For seizure datasets, 18 bipolar channels were first generated from 19 unipolar channels, following [45]. And EEGs were segmented into 4s non-overlapping trials. Then, trials from both datasets were preprocessed by a 50 Hz notch filter and a 0.5-50 Hz bandpass filter. For CHSZ, 1000 Hz trials were downsampled to 500 Hz for consistency. For NICU, the ground truth label of each trial was determined by the majority vote of three experts.
EA was utilized after pre-processing on all datasets, and the EA reference matrix was updated online when new test trials arrived on-the-fly, as in [46].

### v2

@@ Abstract
Electroencephalography (EEG)-based brain-computer interfaces (BCIs) transform spontaneous/evoked neural activity into control commands for external communication. While convolutional neural networks (CNNs) remain the mainstream backbone for EEG decoding, their inherently short receptive field makes it difficult to capture long-range temporal dependencies and global inter-channel relationships. Recent CNN-Transformer (Conformer) hybrids partially address this issue, but most adopt a serial design, resulting in suboptimal integration of local and global features, and often overlook explicit channel-wise modeling. To address these limitations, we propose DBConformer, a dual-branch convolutional Transformer network tailored for EEG decoding. It integrates a temporal Conformer to model long-range temporal dependencies and a spatial Conformer to extract inter-channel interactions, capturing both temporal dynamics and spatial patterns in EEG signals. A lightweight channel attention module further refines spatial representations by assigning data-driven importance to EEG channels. Extensive experiments under four evaluation settings on three paradigms, including motor imagery, seizure detection, and steady-state visual evoked potential, demonstrated that DBConformer consistently outperformed 13 competitive baseline models, with over an eight-fold reduction in parameters than current high-capacity EEG Conformer architecture. Furthermore, the visualization results confirmed that the features extracted by DBConformer are physiologically interpretable and aligned with prior knowledge. The superior performance and interpretability of DBConformer make it reliable for accurate, robust, and explainable EEG decoding. Code is publicized at https://github.com/wzwvv/DBConformer.
@@ I Introduction
A brain-computer interface (BCI) provides a direct communication pathway between cortical activity and external devices [1]. Non-invasive BCIs based on electroencephalography (EEG) have gained prominence because EEG sensors are inexpensive, portable, and safe for extended use [2]. Nevertheless, the low signal-to-noise ratio, non-stationarity, and significant inter-subject variability of scalp EEG impose stringent requirements on the decoding model that translates raw signals into task-relevant information.
This study focuses on three significant EEG decoding tasks:
@@
    Motor imagery (MI) classification [3], which enables brain-based control of devices/prostheses. During MI, users imagine the movement of specific body parts, producing event-related desynchronization/synchronization patterns over the brain’s motor cortex [4].
  2. 2.
    Epileptic seizure detection [5], a task of automated neurological monitoring. Epilepsy, affecting over 70 million people worldwide [6], is characterized by recurrent seizures. Continuous EEG monitoring remains the gold standard for seizure detection.
  3. 3.
    Steady-state visual evoked potential (SSVEP) [7] recognition, which occurs when users fixate on flickering visual stimuli, producing EEG responses at the same stimulus frequency and its harmonics.
The decoding model is essential for effective EEG decoding and has been a research hotspot of BCIs. Three trends have emerged in the design of EEG decoding models:
  1. 1.
    Model Type: Classic convolutional neural networks (CNNs), such as EEGNet [8], Shallow CNN (SCNN), and Deep CNN (DCNN) [9], exploit local spatio-temporal patterns but struggle with global temporal dependencies. CNN-Transformer (Conformer) hybrids extend CNN locality with self-attention based global context. Yet, prior work [10, 11, 12, 13] stacks the two modules serially, restricting interaction between local and global features and neglecting explicit spatial modeling.
@@
Overall, the latest serial Conformer hybrids first compress EEG trials into patch embeddings by CNN blocks and then utilize self-attention to capture global long-range dependencies. Although they are effective for certain tasks, this type of architecture forces all spatial information to pass through a temporal bottleneck, preventing interaction between local and global characteristics. Notably, some Conformer hybrids have large model sizes, e.g., EEG Conformer has around 0.8M parameters, which increases computational cost and may overfit on limited EEG data. Besides, spatial inter-channel relationships, often captured in early CNN layers, may be overwhelmed in later blocks, limiting the model’s capacity to leverage spatial information crucial for MI classification and seizure detection. A balanced, parallel design that treats temporal and spatial patterns symmetrically remains absent.
To address the above limitations, we propose DBConformer, a dual-branch convolutional Transformer architecture that performs parallel spatio-temporal representation learning for EEG decoding. The temporal Conformer (T-Conformer) extracts hierarchical temporal dynamics by integrating depth-wise temporal convolutions with multi-head self-attention, whereas the spatial Conformer (S-Conformer) utilizes point-wise convolutions and channel-wise self-attention to model spatial dependencies. A lightweight channel attention module adaptively refines spatial saliency before branch fusion. In each branch, CNN extracts temporal/spatial local representations, and Transformer captures global dependencies in the temporal or channel horizon. The comparison of network architectures among CNNs, traditional serial Conformers, and the proposed DBConformer, illustrated in Fig. 1.
Figure 1: Comparison of network architectures among CNNs (EEGNet[8], SCNN [9], DCNN [9], etc), traditional serial CNN-Transformer hybrids (EEG Conformer [10], CTNet [11], etc), and the proposed DBConformer. DBConformer has two branches that parallel capture temporal and spatial characteristics.
@@
    Extensive experiments on three paradigms validated that DBConformer almost always outperforms 13 competitive baseline models. Furthermore, it achieves this superior performance with over an eight-fold reduction in parameters than the high-capacity EEG Conformer baseline, demonstrating both effectiveness and efficiency.
@@ II Related Works
FBCNet [14] Parallel 9 Spectral MI
@@
DFformer [32] Serial 1 No MI, sleep stage classification
MI-CAT [33] Serial 1 No MI
MSCFormer [34] Parallel 3 Temporal MI
TMSA-Net [35] Parallel 2 Temporal MI
MSVTNet [36] Parallel 4 Temporal MI
Dual-TSST [37] Parallel 2 Temporal MI, emotion recognition
GAT [38] Serial 1 No MI
@@ II-B Serial CNN-Transformer Hybrids
To incorporate global information, recent EEG decoding architectures incorporate Transformers into CNNs, typically following a serial architecture. Specifically, EEG Conformer [10], CTNet [11], EEG-Deformer [12], SE-TSS-Transformer [13], and DFformer [32] first compress EEG trials into patch embeddings using convolutional blocks, and subsequently apply Transformer layers to capture long-range dependencies. In particular, MI-CAT [33] extracts intra-domain and inter-domain features through a temporal-spatial CNN module followed by stacked Transformer blocks. GAT [38] combines parallel convolutions with an attention adaptor to enhance domain transferability. Similarly, MSCFormer [34] and TMSA-Net [35] integrate multi-branch, multi-scale CNNs with Transformer encoders to jointly model local and global EEG representations. MSVTNet [36] further extracts local spatio-temporal features at different filtered scales and introduces an auxiliary branch loss to facilitate effective CNN-Transformer integration.
Other works employ additional data transformations. For instance, Dual-TSST [37] applies a wavelet transform to EEG signals and then extracts temporal-spatial-spectral representations from both raw and transformed data via dual CNN branches, followed by Transformer-based fusion. MVCNet [17] is a multi-view contrastive network utilizing multi-domain EEG data augmentations and incorporating both cross-view and cross-model modules for MI decoding.
However, this serial design constrains information flow to a single pathway, where local and global features are processed sequentially. As a result, spatial inter-channel relationships, often captured in early CNN layers, may be overwhelmed in later blocks, limiting the model’s capacity to leverage spatial information. Neuroscientific studies further suggest that temporal and spatial information are processed by partly dissociable mechanisms in the brain [39, 40], reinforcing the motivation for architectures that explicitly decouple and then integrate temporal and spatial representations.
@@ II-C Other Architectures
Some research explores architectures beyond CNNs, including full self-attention mechanisms [42, 43], graph neural networks [44, 45, 46, 47], and long short term memory layers [48]. Transformer-based models capture long-range dependencies, and graph neural networks leverage electrode topology to refine spatial characteristics. These models are typically parameter-intensive and rarely integrate local feature extraction with global modeling.
@@ III-B Data Normalization
The raw EEG trials of a batch are $\mathbf{X}\!\in\!\mathbb{R}^{B\times C\times T}$, where $B$ is the batch size, $C$ the number of channels, and $T$ the number of time samples. We apply Euclidean alignment (EA) [49] to align the original trials from each subject in the Euclidean space. Assume a subject has $n$ trials, EA can be formulated as the following for the $i$-th trial:
@@ III-C Network Architecture
T-Conformer Separable Conv1D $F$ $(1,)$ $C\cdot F$ $(B,F,T)$ Spatial filtering
@@
Temporal Temporal Conv1D $F$ $(K,)$ $F\cdot K$ $(B,F,T)$ Temporal depthwise Conv.
@@
III-C1 Motivation of Dual-Branch Design
EEG decoding requires simultaneously modeling fine-grained temporal oscillations and spatially distributed cortical activations. Most existing CNN-Transformer hybrids adopt serial designs, where CNN layers first extract embeddings that are subsequently refined by Transformer blocks. Such designs tend to bias the model toward either local feature extraction (when CNN dominates) or global dependency modeling (when Transformer dominates), leading to limited complementarity.
By contrast, DBConformer explicitly decouples temporal and spatial learning into parallel branches with distinct inductive biases:
  • The temporal branch emphasizes local spectral-temporal patterns while also modeling long-range dependencies.
  • The spatial branch emphasizes inter-channel interactions, adaptively weighted by channel attention.
Formally, a batch of input EEG trials is denoted as $\mathbf{X}\in\mathbb{R}^{B\times C\times T}$, where $B$ is the number of trials, $C$ the number of channels, and $T$ the number of time samples. DBConformer processes $\mathbf{X}$ through two parallel mappings:
$\mathbf{F}_{\text{t}}$ $=f_{\text{T-Conformer}}(\mathbf{X})\in\mathbb{R}^{B\times D},$ (3)
$\mathbf{F}_{\text{s}}$ $=f_{\text{S-Conformer}}(\mathbf{X})\in\mathbb{R}^{B\times D},$ (4)
where $D$ is the embedding dimension, $f_{\text{T-Conformer}}(\cdot)$ captures temporal dependencies, and $f_{\text{S-Conformer}}(\cdot)$ captures spatial relations.
This parallel design ensures that temporal and spatial features are learned independently before integration, yielding more disentangled and complementary representations. Unlike prior serial hybrids, this design is further supported by neuroscientific evidence. For example, neurophysiological studies demonstrated that absolute duration and distance are independently coded by distinct populations of prefrontal neurons [39], whereas behavioral studies showed that temporal and spatial attention recruit separable neural systems [40]. Such findings suggest that temporal rhythms and spatial topographies are processed by distinct but interacting cortical mechanisms, providing both physiological plausibility and theoretical grounding for our dual-branch architecture.
III-C2 T-Conformer
@@
Inspired by prior CNN backbones in EEG decoding [15], we design the module with a depthwise separable 1D convolution, a 1D temporal convolution, and an average pooling layer. The first separable convolution has $F$ kernels with a stride of $(1,)$, followed with batch normalization. The second temporal convolution applies $F$ filters, accompanied by batch normalization, a GELU activation, and a dropout with a probability $p=0.5$. Then, an average pooling with kernel size $(1,W)$ is applied along the time dimension to form non-overlapping temporal patches $\mathbf{Z}_{\text{t}}$. Finally, the patches are reshaped via transposition. The number of filters $F$ is set equal to the Transformer embedding dimension $D$, thereby eliminating the need for a separate linear projection.
@@
$\mathbf{F}_{\text{t}}=\frac{1}{P}\sum_{i=1}^{P}\mathbf{Z}_{t}^{\text{out}}[\mathrel{\mathop{\ordinarycolon}},i,\mathrel{\mathop{\ordinarycolon}}]\in\mathbb{R}^{B\times D},$ (8)
@@ III-C3 S-Conformer
Figure 3: The proposed channel attention module.
@@
Then, for each sample, the $c$-th channel token $\mathbf{z}_{i}^{c}$ is aggregated through an attention-weighted sum:
@@
This operation is executed independently for each sample, yielding its spatial representation $\mathbf{f}_{\text{s},i}\in\mathbb{R}^{D}$. By repeating this computation across all $B$ samples in a batch, the batch-level spatial representation $\mathbf{F}_{\text{s}}\in\mathbb{R}^{B\times D}$ can be obtained.
III-C4 Feature Fusion and Classification
@@
which is then passed through a multi-layer perceptron classifier to generate the final predictions $\hat{y}=\mathrm{softmax}\big(\mathrm{MLP}(\mathbf{F}_{\text{fused}})\big)$. The classifier consists of three fully connected layers with intermediate ELU activations and dropout regularization.
@@ IV Experiments and Results
This section details the datasets, experiments, and analyses. We implemented and fairly evaluated thirteen state-of-the-art EEG decoding models, including CNN-based models, CNN-Transformer hybrids, and CNN-Mamba hybrids. Code for DBConformer, along with all compared baseline models, is available on GitHub [^1] https://github.com/wzwvv/DBConformer, serving as a benchmark codebase for EEG decoding.
@@ IV-A1 MI Datasets
We adopted six publicly available MI datasets from the MOABB benchmark [51] and BCI competitions:
@@
  • Zhou2016 [54]: Contains EEG trials from 4 subjects performing left/right hand MI, recorded with 14 channels at 250 Hz. The first session was used in the experiments.
  • Blankertz2007 [55]: Contains EEG trials from 7 subjects with 59 channels EEG sampled at 250 Hz from BCI Competition IV dataset 1. Subjects 1 and 5 performed left hand/right foot MI, and the other five subjects performed left/right hand MI.
  • BNCI2014002 [56]: Contains EEG trials from 14 subjects, each performing eight runs right hand/feet MI, using 15 EEG channels recorded at 512 Hz. The first five training runs were used in the experiments.
  • OpenBMI [57]: Contains EEG trials from 54 subjects, each performing two sessions of left/right hand MI, using 62 channels recorded at 1,000 Hz. The total number of trials is 10,800.
@@ IV-A2 Seizure Detection Datasets
IV-A3 SSVEP Dataset
The Nakanishi2015 [60] SSVEP dataset is evaluated. Data were recorded at 256 Hz from 10 subjects, with eight channels PO7, PO3, POZ, PO4, PO8, O1, Oz and O2 in the occipital area. For each subject, 12-class stimuli were recorded with 15 blocks.
TABLE III: Summary of the six MI datasets, two seizure datasets, and one SSVEP dataset.
@@ IV-A3 SSVEP Dataset
SSVEP Nakanishi2015 10 8 256 1 1,800 12 different stimuli
IV-A4 Data Preprocessing
For the MI paradigm, the standard preprocessing steps in MOABB, including notch filtering, band-pass filtering, etc., were performed to ensure reproducibility on BNCI2014001, BNCI2014004, Zhou2016, and BNCI2014002 datasets. For the Blankertz2007 dataset, EEG trials were first band-pass filtered between 8 and 30 Hz. Then, trials between [0.5, 3.5] seconds after the cue onset were segmented and downsampled to 250 Hz. For the OpenBMI dataset, the raw EEG signals were first filtered by a third-order Butterworth band-pass filter with a passband of [4, 40] Hz to remove low-frequency drifts and high-frequency noise, and then downsampled to 256 Hz. Following previous studies such as FBCNet [14] and FBMSNet [21], 20 motor-related channels {FC5, FC3, FC1, FC2, FC4, FC6, C5, C3, C1, Cz, C2, C4, C6, CP5, CP3, CP1, CPz, CP2, CP4, CP6} were selected for analysis, which cover the sensorimotor cortex regions typically associated with MI tasks. EA was utilized after pre-processing on all MI datasets, and the EA reference matrix was updated online when new test trials arrived on-the-fly, as in [61].
For seizure datasets, 18 bipolar channels were first generated from 19 unipolar channels, following [59]. EEG recordings were segmented into 4s non-overlapping trials. Then, trials from both datasets were preprocessed by a 50 Hz notch filter and a 0.5-50 Hz bandpass filter. For CHSZ, 1000 Hz trials were downsampled to 500 Hz for consistency. For NICU, the ground-truth label of each trial was determined by the majority vote of three experts.
For the SSVEP paradigm, EEG trials were first down-sampled to 256 Hz and [6, 80] Hz band-pass filtered, and then split into 1-second length trials in Nakanishi2015, following the preprocessing steps in [62].